# Improving Diffusion Inverse Problem Solving with Decoupled Noise Annealing

### v2

@@ 2.2 Bayesian Inverse Problems with Diffusion
$-2\dot{\sigma}_{t}\sigma_{t}\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{y}}\mid{\mathbf{x}}_{t})\mathrm{d}t+\sqrt{2\dot{\sigma}_{t}\sigma_{t}}\mathrm{d}{\mathbf{w}}_{t}.$ (1)
Here, the noisy likelihood $\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{y}}\mid{\mathbf{x}}_{t})$ is generally intractable. Multiple methods have been proposed to estimate the noisy likelihood [50, 12, 46, 5, 36, 54, 8, 59, 41]. One predominant approach is the DPS algorithm [13], which estimates $p({\mathbf{y}}\mid{\mathbf{x}}_{t})\approx p({\mathbf{y}}\mid{\mathbf{x}}_{0}=\mathbb{E}[{\mathbf{x}}_{0}\mid{\mathbf{x}}_{t}])$. Another line of work [30, 29, 54, 7, 18] solves linear inverse problems by running the reverse diffusion in the spectral domain via singular value decomposition (SVD). Other methods bypass direct computation of this likelihood by interleaving optimization [45, 24, 9, 2, 59, 31, 56] or projection [30, 10, 23, 54] steps with normal diffusion sampling steps.
@@ 3.2 Practical Design Choice of DAPS
$\approx\nabla_{{\mathbf{x}}_{0}}\log p({\mathbf{x}}_{t}\mid{\mathbf{x}}_{0})+{\bm{s}}_{\bm{\theta}}({\mathbf{x}}_{0},t_{\min}).$ (6)
@@ 3.4 Discussion with Existing Methods
Comparison with other posterior sampling methods. Unlike most previous algorithms, our sampling method does not solve a specific SDE/ODE; instead, we recursively sample from the time-marginal $p({\mathbf{x}}_{t}\mid{\mathbf{y}})$ with noise annealing to zero. We therefore decouple the dependency of ${\mathbf{x}}_{t}$ and ${\mathbf{x}}_{t+\Delta t}$ in the sampling process. We argue that decoupling helps correct the errors accumulated in early diffusion sampling steps by allowing non-local transitions. This is particularly important when the measurement function is nonlinear. As a concrete example, Fig. 4 compares DAPS and DPS when solving a 2D nonlinear inverse problem with a Gaussian mixture prior. We visualize the trajectories of ${\mathbf{x}}_{t}$ from time $T$ to 00 for both methods. For DAPS, points on the trajectory have significantly larger variations compared to those of DPS. As a result, DPS converges to wrong solutions, but DAPS is able to approximate the true posterior distribution. We provide more discussions in Appendix G.
@@ 4.1 Experimental Setup
As mentioned in Sec. 3, we implement a few-step Euler ODE solver to compute $\hat{\mathbf{x}}_{0}({\mathbf{x}}_{t})$, while maintaining the same number of neural function evaluations (NFE) for different noise levels. In our experiments, we use DAPS-1K for all linear tasks and DAPS-4K for all nonlinear tasks. DAPS-1K uses $4$ ODE solver NFE and $250$ annealing scheduling steps, while DAPS-4K uses $10$ NFE and $400$ steps, respectively. We discuss the effect of choosing these configurations in Sec. 4.4. The corresponding latent version, LatentDAPS, follows the same settings but performs sampling in the latent space of VAEs. We use $100$ and $50$ Langevin steps per denoising iteration for DAPS and LatentDAPS respectively, and tune learning rates separately for each task. Further details on model configurations, samplers, and other hyperparameters are provided in Appendix F, along with more discussion on sampling efficiency in Fig. 6(b) and Sec. E.1.
@@
Inverse problems. We evaluate our method with a series of linear and nonlinear tasks. For linear inverse problems, we consider (1) super-resolution, (2) Gaussian deblurring, (3) motion deblurring, (4) inpainting (with a box mask), and (5) inpainting (with a 70% random mask). For Gaussian and motion deblurring, kernels of size 61$\times$61 with standard deviations of 3.0 and 0.5, respectively, are used. In the super-resolution task, a bicubic resizer downscales images by a factor of 4. The box inpainting task uses a random box of size 128$\times$128 to mask the original images, while random mask inpainting uses a generated random mask where each pixel has a 70% chance of being masked, following the settings in [45].
@@
Baselines. We compare our methods with the following baselines: DDRM [30], DPS [13], DDNM [54], DCDP [31], FPS-SMC [18], DiffPIR [59] for pixel-based diffusion model experiments. We compare the latent diffusion version of our methods with PSLD [42] and ReSample [45], which also operate in the latent space. Note that DDRM, DDNM, FPS-SMC, and PSLD cannot handle nonlinear inverse problems. We especially RED-diff [35] for nonlinear experiments and a traditional method HIO [14] for phase retrieval.
@@
Task Type Method FFHQ ImageNet PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) FID ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) FID ($\downarrow$) Super Resolution 4× Pixel DAPS (ours) 29.07 0.818 0.177 $\underline{51.44}$ 25.89 $\underline{0.694}$ 0.276 83.57 DPS $25.86$ $0.753$ $0.269$ $81.07$ $21.13$ $0.489$ $0.361$ $106.32$ DDRM $26.58$ $0.782$ $0.282$ $79.25$ $22.62$ $0.521$ $0.324$ $103.85$ DDNM $28.03$ $0.795$ $0.197$ $64.62$ $23.96$ $0.604$ $0.475$ $98.62$ DCDP $\underline{28.66}$ $0.807$ $\underline{0.178}$ $53.81$ FPS-SMC $28.42$ $\underline{0.813}$ $0.204$ 49.25 $\underline{24.82}$ 0.703 $\underline{0.313}$ $\underline{97.51}$ DiffPIR $26.64$ $0.260$ $65.77$ $23.18$ $0.371$ $106.32$ Latent LatentDAPS(ours) 27.48 0.801 0.182 $59.62$ $\underline{25.06}$ $\underline{0.673}$ 0.276 84.37 PSLD $\underline{24.35}$ $\underline{0.649}$ $\underline{0.287}$ $74.36$ 25.42 0.694 $\underline{0.360}$ $\underline{97.45}$ ReSample $23.29$ $0.594$ $0.392$ $93.18$ $22.61$ $0.576$ $0.370$ $113.42$ Inpaint (Box) Pixel DAPS(ours) $24.07$ $0.814$ 0.133 43.10 $21.43$ $0.725$ $\underline{0.214}$ $\underline{109.85}$ DPS $22.51$ $0.792$ $0.209$ $61.27$ $18.94$ $0.722$ $0.257$ $126.52$ DDRM $22.26$ $0.801$ $0.207$ $78.62$ $18.63$ $0.733$ $0.254$ $116.37$ DDNM $\underline{24.47}$ 0.837 $0.235$ $46.59$ $\underline{21.64}$ 0.748 $0.319$ 103.97 DCDP $23.89$ $0.760$ $0.163$ $\underline{45.23}$ FPS-SMC 24.86 $\underline{0.823}$ $\underline{0.146}$ $48.34$ 22.16 $\underline{0.726}$ 0.208 $111.58$ Latent LatentDAPS(ours) $\underline{23.99}$ $\underline{0.802}$ $0.194$ $\underline{46.52}$ $17.19$ $0.624$ $\underline{0.340}$ $\underline{145.63}$ PSLD 24.22 0.813 0.158 43.02 20.10 0.694 $0.465$ $146.53$ ReSample $20.06$ $0.749$ $\underline{0.184}$ $53.21$ $\underline{18.29}$ $\underline{0.631}$ 0.262 127.84 Inpaint (Random) Pixel DAPS(ours) 31.12 0.844 0.098 32.17 $\underline{28.44}$ $\underline{0.775}$ 0.135 54.25 DPS $25.46$ $0.823$ $0.203$ $69.20$ $23.52$ $0.745$ $0.297$ $87.53$ DDNM $29.91$ $0.817$ $\underline{0.121}$ $\underline{44.37}$ 31.16 0.841 $\underline{0.191}$ $\underline{63.84}$ DCDP $\underline{30.69}$ $\underline{0.842}$ $0.142$ $52.51$ FPS-SMC $28.21$ $0.823$ $0.261$ $61.23$ $24.52$ $0.701$ $0.316$ $79.12$ Latent LatentDAPS(ours) 30.71 0.813 $\underline{0.141}$ 36.41 $\underline{27.59}$ $\underline{0.772}$ $\underline{0.164}$ $\underline{61.62}$ PSLD $\underline{30.31}$ $\underline{0.809}$ $0.221$ $47.21$ 31.30 0.783 $0.337$ $83.21$ ReSample $29.61$ $0.746$ 0.140 $\underline{39.85}$ $27.50$ $0.756$ 0.143 59.87 Gaussian deblurring Pixel DAPS(ours) 29.19 0.817 0.165 53.33 $\underline{26.15}$ $\underline{0.684}$ 0.253 75.68 DPS $25.87$ $0.764$ $0.219$ $79.75$ $20.31$ $0.598$ $0.397$ $116.42$ DDRM $24.93$ $0.732$ $0.239$ $92.43$ $21.26$ $0.564$ $0.443$ $146.89$ DDNM $\underline{28.20}$ $\underline{0.804}$ $\underline{0.216}$ $\underline{57.83}$ 28.06 0.703 $\underline{0.278}$ $\underline{81.43}$ DCDP $27.50$ $0.699$ $0.304$ $86.43$ FPS-SMC $26.54$ $0.773$ $0.253$ $67.45$ $23.91$ $0.601$ $0.387$ $91.72$ DiffPIR $27.36$ $0.236$ $59.65$ $22.80$ $0.355$ $93.36$ Latent LatentDAPS(ours) 27.93 0.764 0.234 64.52 $25.05$ $0.668$ $\underline{0.345}$ $\underline{78.51}$ PSLD $23.27$ $0.631$ $0.316$ $89.51$ $\underline{25.86}$ $\underline{0.688}$ $0.390$ $91.39$ ReSample $\underline{26.39}$ $\underline{0.714}$ $\underline{0.255}$ $\underline{71.69}$ 25.97 0.703 0.254 65.35 Motion deblurring Pixel DAPS(ours) 29.66 0.847 0.157 39.49 27.86 0.766 0.196 61.83 DPS $24.52$ $0.801$ $0.246$ $65.23$ $18.96$ $0.629$ $0.423$ $137.81$ DCDP $25.08$ $0.512$ $0.364$ $125.13$ FPS-SMC $\underline{27.39}$ $\underline{0.826}$ $\underline{0.227}$ $\underline{48.32}$ $\underline{24.52}$ $\underline{0.647}$ $\underline{0.326}$ $\underline{87.43}$ DiffPIR $26.57$ $0.255$ $65.78$ $24.01$ $0.366$ $94.63$ Latent LatentDAPS(ours) $\underline{27.00}$ $\underline{0.814}$ $\underline{0.283}$ $\underline{46.84}$ $\underline{26.83}$ 0.745 $\underline{0.296}$ $\underline{73.62}$ PSLD $22.31$ $0.678$ $0.336$ $96.15$ $20.85$ $0.594$ $0.511$ $124.67$ ReSample 27.41 0.823 0.198 44.72 26.94 $\underline{0.738}$ 0.227 66.89 Phase Retrieval Pixel DAPS(ours) $\textbf{30.63}_{\pm 3.13}$ $\textbf{0.851}_{\pm 0.072}$ $\textbf{0.139}_{\pm 0.060}$ 42.71 $\textbf{25.78}_{\pm 6.92}$ $\textbf{0.743}_{\pm 0.084}$ $\textbf{0.254}_{\pm 0.125}$ 82.67 DPS $17.64_{\pm 2.97}$ $0.441_{\pm 0.129}$ $0.410_{\pm 0.090}$ $104.52$ $\underline{16.81}_{\pm 3.61}$ $\underline{0.427}_{\pm 0.143}$ $\underline{0.447}_{\pm 0.099}$ $\underline{197.54}$ RED-diff $15.60_{\pm 4.48}$ $0.398_{\pm 0.195}$ $0.596_{\pm 0.092}$ $167.43$ $14.98_{\pm 3.75}$ $0.386_{\pm 0.057}$ $0.536_{\pm 0.129}$ $212.24$ DCDP $\underline{28.65}_{\pm 8.09}$ $\underline{0.781}_{\pm 0.217}$ $\underline{0.203}_{\pm 0.196}$ $\underline{68.13}$ Latent LatentDAPS(ours) $\textbf{29.16}_{\pm 3.55}$ $\textbf{0.796}_{\pm 0.089}$ $\textbf{0.199}_{\pm 0.078}$ 54.26 $\textbf{20.54}_{\pm 6.41}$ $\textbf{0.612}_{\pm 0.114}$ $\textbf{0.361}_{\pm 0.150}$ 129.54 ReSample $21.60_{\pm 8.10}$ $0.648_{\pm 0.154}$ $0.406_{\pm 0.224}$ $84.32$ $19.24_{\pm 4.21}$ $0.618_{\pm 0.146}$ $0.403_{\pm 0.174}$ $130.47$ Classical HIO $13.53_{\pm 2.50}$ $0.359_{\pm 0.093}$ $0.726_{\pm 0.068}$ $268.09$ Nonlinear deblur Pixel DAPS(ours) $\underline{28.29}_{\pm 1.77}$ $\underline{0.783}_{\pm 0.036}$ $\textbf{0.155}_{\pm 0.032}$ $\underline{49.38}$ $\underline{27.73}_{\pm 3.23}$ $\underline{0.724}_{\pm 0.048}$ $\textbf{0.169}_{\pm 0.056}$ $\underline{59.87}$ DPS $23.39_{\pm 2.01}$ $0.623_{\pm 0.082}$ $0.278_{\pm 0.060}$ $91.31$ $22.49_{\pm 3.20}$ $0.591_{\pm 0.101}$ $0.306_{\pm 0.081}$ $101.41$ RED-diff $\textbf{30.86}_{\pm 0.51}$ $\textbf{0.795}_{\pm 0.028}$ $\underline{0.160}_{\pm 0.034}$ 43.84 $\textbf{30.07}_{\pm 1.41}$ $\textbf{0.754}_{\pm 0.023}$ $\underline{0.211}_{\pm 0.083}$ 51.22 DCDP $27.92_{\pm 2.64}$ $0.779_{\pm 0.067}$ $0.183_{\pm 0.051}$ $51.96$ Latent LatentDAPS(ours) $28.11_{\pm 1.75}$ $0.713_{\pm 0.041}$ $0.235_{\pm 0.049}$ $53.63$ $25.34_{\pm 3.44}$ $0.615_{\pm 0.057}$ $0.314_{\pm 0.080}$ $76.73$ ReSample $\textbf{28.24}_{\pm 1.69}$ $0.742_{\pm 0.039}$ $\textbf{0.185}_{\pm 0.039}$ 51.62 $\textbf{26.20}_{\pm 3.71}$ $\textbf{0.653}_{\pm 0.064}$ $\textbf{0.206}_{\pm 0.057}$ 61.16 High dynamic range Pixel DAPS(ours) $\textbf{27.12}_{\pm 3.53}$ $\textbf{0.752}_{\pm 0.041}$ $\textbf{0.162}_{\pm 0.072}$ 42.97 $\textbf{26.30}_{\pm 4.10}$ $\textbf{0.717}_{\pm 0.067}$ $\textbf{0.175}_{\pm 0.107}$ 64.19 DPS $\underline{22.73}_{\pm 6.07}$ $\underline{0.591}_{\pm 0.141}$ $0.264_{\pm 0.156}$ $112.82$ $19.23_{\pm 2.52}$ $0.582_{\pm 0.082}$ $0.503_{\pm 0.106}$ $146.23$ RED-diff $22.16_{\pm 3.41}$ $0.512_{\pm 0.083}$ $\underline{0.258}_{\pm 0.089}$ $\underline{108.32}$ $\underline{22.03}_{\pm 5.90}$ $\underline{0.601}_{\pm 0.094}$ $\underline{0.274}_{\pm 0.198}$ $\underline{113.48}$ Latent LatentDAPS(ours) $\textbf{25.94}_{\pm 2.87}$ $\textbf{0.751}_{\pm 0.056}$ $0.223_{\pm 0.080}$ $74.83$ $23.64_{\pm 4.10}$ $0.609_{\pm 0.053}$ $0.269_{\pm 0.099}$ $93.51$ ReSample $25.65_{\pm 3.57}$ $0.732_{\pm 0.059}$ $\textbf{0.182}_{\pm 0.085}$ 67.22 $\textbf{25.11}_{\pm 4.21}$ $\textbf{0.633}_{\pm 0.049}$ $\textbf{0.198}_{\pm 0.089}$ 87.66

### v3

@@ 2.2 Bayesian Inverse Problems with Diffusion
$\hskip-10.0pt-2\dot{\sigma}_{t}\sigma_{t}\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{y}}\mid{\mathbf{x}}_{t})\mathrm{d}t+\sqrt{2\dot{\sigma}_{t}\sigma_{t}}\mathrm{d}{\mathbf{w}}_{t}.$ (1)
Here, the noisy likelihood $\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{y}}\mid{\mathbf{x}}_{t})$ is generally intractable. Multiple methods have been proposed to estimate the noisy likelihood [50, 12, 46, 5, 36, 54, 8, 60, 41]. One predominant approach is the DPS algorithm [13], which estimates $p({\mathbf{y}}\mid{\mathbf{x}}_{t})\approx p({\mathbf{y}}\mid{\mathbf{x}}_{0}=\mathbb{E}[{\mathbf{x}}_{0}\mid{\mathbf{x}}_{t}])$. Another line of work [30, 29, 54, 7, 18] solves linear inverse problems by running the reverse diffusion in the spectral domain via singular value decomposition (SVD). Other methods bypass direct computation of this likelihood by interleaving optimization [45, 24, 9, 2, 60, 31, 56, 57] or projection [30, 10, 23, 54] steps with normal diffusion sampling steps.
@@ 3.2 Practical Design Choice of DAPS
$\hskip-20.00003pt\approx\nabla_{{\mathbf{x}}_{0}}\log p({\mathbf{x}}_{t}\mid{\mathbf{x}}_{0})+{\bm{s}}_{\bm{\theta}}({\mathbf{x}}_{0},t_{\min}).$ (6)
@@ 3.4 Discussion with Existing Methods
Comparison with other posterior sampling methods. Unlike most previous algorithms, our sampling method does not solve a specific SDE/ODE; instead, we recursively sample from the time-marginal $p({\mathbf{x}}_{t}\mid{\mathbf{y}})$ with noise annealing to zero. We therefore decouple the dependency of ${\mathbf{x}}_{t}$ and ${\mathbf{x}}_{t+\Delta t}$ in the sampling process. We argue that decoupling helps correct the errors accumulated in early diffusion sampling steps by allowing non-local transitions. This is particularly important when the measurement function is nonlinear. As a concrete example, Fig. 4 compares DAPS and DPS when solving a 2D nonlinear inverse problem with a Gaussian mixture prior. We visualize the trajectories of ${\mathbf{x}}_{t}$ from time $T$ to 0 for both methods. For DAPS, points on the trajectory have significantly larger variations compared to those of DPS. As a result, DPS converges to wrong solutions, but DAPS is able to approximate the true posterior distribution. We provide more discussions in Appendix G.
@@ 4.1 Experimental Setup
As mentioned in Sec. 3, we implement a few-step Euler ODE solver to compute $\hat{\mathbf{x}}_{0}({\mathbf{x}}_{t})$, while maintaining the same number of neural function evaluations (NFE) for different noise levels. In our experiments, we use DAPS-1K for all linear tasks and DAPS-4K for all nonlinear tasks. DAPS-1K uses $4$ ODE solver NFE and $250$ annealing scheduling steps, while DAPS-4K uses $10$ NFE and $400$ steps, respectively. We discuss the effect of choosing these configurations in Sec. 4.4. The corresponding latent version, LatentDAPS, follows the same settings but performs sampling in the latent space of VAEs. We use $100$ and $50$ Langevin steps per denoising iteration for DAPS and LatentDAPS respectively, and tune learning rates separately for each task. Further details on model configurations, samplers, and other hyperparameters are provided in Appendix F, along with more discussion on sampling efficiency in Fig. 6 and Sec. E.1.
@@
Inverse problems. We evaluate our method with a series of linear and nonlinear tasks. For linear inverse problems, we consider (1) super-resolution, (2) Gaussian deblurring, (3) motion deblurring, (4) inpainting (with a box mask), and (5) inpainting (with a 70% random mask). For Gaussian and motion deblurring, kernels of size $\times61$ with standard deviations of 3.0 and 0.5, respectively, are used. In the super-resolution task, a bicubic resizer downscales images by a factor of 4. The box inpainting task uses a random box of size $\times128$ to mask the original images, while random mask inpainting uses a generated random mask where each pixel has a 70% chance of being masked, following the settings in [45].
@@
Baselines. We compare our methods with the following baselines: DDRM [30], DPS [13], DDNM [54], DCDP [31], FPS-SMC [18], DiffPIR [60], DPnP [57] for pixel-based diffusion model experiments. We compare the latent diffusion version of our methods with PSLD [42] and ReSample [45], which also operate in the latent space. Note that DDRM, DDNM, FPS-SMC, and PSLD cannot handle nonlinear inverse problems. We especially RED-diff [35] for nonlinear experiments and a traditional method HIO [14] for phase retrieval.
@@
Task Type Method FFHQ ImageNet PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) FID ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) FID ($\downarrow$) Super resolution 4× Pixel DAPS (ours) 29.07 0.818 0.177 $\underline{51.44}$ 25.89 $\underline{0.694}$ 0.276 83.57 DPS $25.86$ $0.753$ $0.269$ $81.07$ $21.13$ $0.489$ $0.361$ $106.32$ DDRM $26.58$ $0.782$ $0.282$ $79.25$ $22.62$ $0.521$ $0.324$ $103.85$ DDNM $28.03$ $0.795$ $0.197$ $64.62$ $23.96$ $0.604$ $0.475$ $98.62$ DCDP $\underline{28.66}$ $0.807$ $\underline{0.178}$ $53.81$ - - - - FPS-SMC $28.42$ $\underline{0.813}$ $0.204$ 49.25 $\underline{24.82}$ 0.703 $\underline{0.313}$ $\underline{97.51}$ DiffPIR $26.64$ - $0.260$ $65.77$ $23.18$ - $0.371$ $106.32$ Latent LatentDAPS(ours) 27.48 0.801 0.182 $59.62$ $\underline{25.06}$ $\underline{0.673}$ 0.276 84.37 PSLD $\underline{24.35}$ $\underline{0.649}$ $\underline{0.287}$ $74.36$ 25.42 0.694 $\underline{0.360}$ $\underline{97.45}$ ReSample $23.29$ $0.594$ $0.392$ $93.18$ $22.61$ $0.576$ $0.370$ $113.42$ Inpaint (box) Pixel DAPS(ours) $24.07$ $0.814$ 0.133 43.10 $21.43$ $0.725$ $\underline{0.214}$ $\underline{109.85}$ DPS $22.51$ $0.792$ $0.209$ $61.27$ $18.94$ $0.722$ $0.257$ $126.52$ DDRM $22.26$ $0.801$ $0.207$ $78.62$ $18.63$ $0.733$ $0.254$ $116.37$ DDNM $\underline{24.47}$ 0.837 $0.235$ $46.59$ $\underline{21.64}$ 0.748 $0.319$ 103.97 DCDP $23.89$ $0.760$ $0.163$ $\underline{45.23}$ - - - - FPS-SMC 24.86 $\underline{0.823}$ $\underline{0.146}$ $48.34$ 22.16 $\underline{0.726}$ 0.208 $111.58$ Latent LatentDAPS(ours) $\underline{23.99}$ $\underline{0.802}$ $0.194$ $\underline{46.52}$ $17.19$ $0.624$ $\underline{0.340}$ $\underline{145.63}$ PSLD 24.22 0.813 0.158 43.02 20.10 0.694 $0.465$ $146.53$ ReSample $20.06$ $0.749$ $\underline{0.184}$ $53.21$ $\underline{18.29}$ $\underline{0.631}$ 0.262 127.84 Inpaint (random) Pixel DAPS(ours) 31.12 0.844 0.098 32.17 $\underline{28.44}$ $\underline{0.775}$ 0.135 54.25 DPS $25.46$ $0.823$ $0.203$ $69.20$ $23.52$ $0.745$ $0.297$ $87.53$ DDNM $29.91$ $0.817$ $\underline{0.121}$ $\underline{44.37}$ 31.16 0.841 $\underline{0.191}$ $\underline{63.84}$ DCDP $\underline{30.69}$ $\underline{0.842}$ $0.142$ $52.51$ - - - - FPS-SMC $28.21$ $0.823$ $0.261$ $61.23$ $24.52$ $0.701$ $0.316$ $79.12$ Latent LatentDAPS(ours) 30.71 0.813 $\underline{0.141}$ 36.41 $\underline{27.59}$ $\underline{0.772}$ $\underline{0.164}$ $\underline{61.62}$ PSLD $\underline{30.31}$ $\underline{0.809}$ $0.221$ $47.21$ 31.30 0.783 $0.337$ $83.21$ ReSample $29.61$ $0.746$ 0.140 $\underline{39.85}$ $27.50$ $0.756$ 0.143 59.87 Gaussian deblurring Pixel DAPS(ours) 29.19 0.817 0.165 53.33 $\underline{26.15}$ $\underline{0.684}$ 0.253 75.68 DPS $25.87$ $0.764$ $0.219$ $79.75$ $20.31$ $0.598$ $0.397$ $116.42$ DDRM $24.93$ $0.732$ $0.239$ $92.43$ $21.26$ $0.564$ $0.443$ $146.89$ DDNM $\underline{28.20}$ $\underline{0.804}$ $\underline{0.216}$ $\underline{57.83}$ 28.06 0.703 $\underline{0.278}$ $\underline{81.43}$ DCDP $27.50$ $0.699$ $0.304$ $86.43$ - - - - FPS-SMC $26.54$ $0.773$ $0.253$ $67.45$ $23.91$ $0.601$ $0.387$ $91.72$ DiffPIR $27.36$ - $0.236$ $59.65$ $22.80$ - $0.355$ $93.36$ Latent LatentDAPS(ours) 27.93 0.764 0.234 64.52 $25.05$ $0.668$ $\underline{0.345}$ $\underline{78.51}$ PSLD $23.27$ $0.631$ $0.316$ $89.51$ $\underline{25.86}$ $\underline{0.688}$ $0.390$ $91.39$ ReSample $\underline{26.39}$ $\underline{0.714}$ $\underline{0.255}$ $\underline{71.69}$ 25.97 0.703 0.254 65.35 Motion deblurring Pixel DAPS(ours) 29.66 0.847 0.157 39.49 27.86 0.766 0.196 61.83 DPS $24.52$ $0.801$ $0.246$ $65.23$ $18.96$ $0.629$ $0.423$ $137.81$ DCDP $25.08$ $0.512$ $0.364$ $125.13$ - - - - FPS-SMC $\underline{27.39}$ $\underline{0.826}$ $\underline{0.227}$ $\underline{48.32}$ $\underline{24.52}$ $\underline{0.647}$ $\underline{0.326}$ $\underline{87.43}$ DiffPIR $26.57$ - $0.255$ $65.78$ $24.01$ - $0.366$ $94.63$ Latent LatentDAPS(ours) $\underline{27.00}$ $\underline{0.814}$ $\underline{0.283}$ $\underline{46.84}$ $\underline{26.83}$ 0.745 $\underline{0.296}$ $\underline{73.62}$ PSLD $22.31$ $0.678$ $0.336$ $96.15$ $20.85$ $0.594$ $0.511$ $124.67$ ReSample 27.41 0.823 0.198 44.72 26.94 $\underline{0.738}$ 0.227 66.89 Phase retrieval Pixel DAPS(ours) $\textbf{30.63}_{\pm 3.13}$ $\textbf{0.851}_{\pm 0.072}$ $\textbf{0.139}_{\pm 0.060}$ 42.71 $\textbf{25.78}_{\pm 6.92}$ $\textbf{0.743}_{\pm 0.084}$ $\textbf{0.254}_{\pm 0.125}$ 82.67 DPS $17.64_{\pm 2.97}$ $0.441_{\pm 0.129}$ $0.410_{\pm 0.090}$ $104.52$ $\underline{16.81}_{\pm 3.61}$ $\underline{0.427}_{\pm 0.143}$ $\underline{0.447}_{\pm 0.099}$ $\underline{197.54}$ RED-diff $15.60_{\pm 4.48}$ $0.398_{\pm 0.195}$ $0.596_{\pm 0.092}$ $167.43$ $14.98_{\pm 3.75}$ $0.386_{\pm 0.057}$ $0.536_{\pm 0.129}$ $212.24$ DCDP $\underline{28.65}_{\pm 8.09}$ $\underline{0.781}_{\pm 0.217}$ $\underline{0.203}_{\pm 0.196}$ $\underline{68.13}$ - - - - Latent LatentDAPS(ours) $\textbf{29.16}_{\pm 3.55}$ $\textbf{0.796}_{\pm 0.089}$ $\textbf{0.199}_{\pm 0.078}$ 54.26 $\textbf{20.54}_{\pm 6.41}$ $\textbf{0.612}_{\pm 0.114}$ $\textbf{0.361}_{\pm 0.150}$ 129.54 ReSample $21.60_{\pm 8.10}$ $0.648_{\pm 0.154}$ $0.406_{\pm 0.224}$ $84.32$ $19.24_{\pm 4.21}$ $0.618_{\pm 0.146}$ $0.403_{\pm 0.174}$ $130.47$ Classical HIO $13.53_{\pm 2.50}$ $0.359_{\pm 0.093}$ $0.726_{\pm 0.068}$ $268.09$ - - - - Nonlinear deblur Pixel DAPS(ours) $\underline{28.29}_{\pm 1.77}$ $\underline{0.783}_{\pm 0.036}$ $\textbf{0.155}_{\pm 0.032}$ $\underline{49.38}$ $\underline{27.73}_{\pm 3.23}$ $\underline{0.724}_{\pm 0.048}$ $\textbf{0.169}_{\pm 0.056}$ $\underline{59.87}$ DPS $23.39_{\pm 2.01}$ $0.623_{\pm 0.082}$ $0.278_{\pm 0.060}$ $91.31$ $22.49_{\pm 3.20}$ $0.591_{\pm 0.101}$ $0.306_{\pm 0.081}$ $101.41$ RED-diff $\textbf{30.86}_{\pm 0.51}$ $\textbf{0.795}_{\pm 0.028}$ $\underline{0.160}_{\pm 0.034}$ 43.84 $\textbf{30.07}_{\pm 1.41}$ $\textbf{0.754}_{\pm 0.023}$ $\underline{0.211}_{\pm 0.083}$ 51.22 DCDP $27.92_{\pm 2.64}$ $0.779_{\pm 0.067}$ $0.183_{\pm 0.051}$ $51.96$ - - - - Latent LatentDAPS(ours) $28.11_{\pm 1.75}$ $0.713_{\pm 0.041}$ $0.235_{\pm 0.049}$ $53.63$ $25.34_{\pm 3.44}$ $0.615_{\pm 0.057}$ $0.314_{\pm 0.080}$ $76.73$ ReSample $\textbf{28.24}_{\pm 1.69}$ $0.742_{\pm 0.039}$ $\textbf{0.185}_{\pm 0.039}$ 51.62 $\textbf{26.20}_{\pm 3.71}$ $\textbf{0.653}_{\pm 0.064}$ $\textbf{0.206}_{\pm 0.057}$ 61.16 High dynamic range Pixel DAPS(ours) $\textbf{27.12}_{\pm 3.53}$ $\textbf{0.752}_{\pm 0.041}$ $\textbf{0.162}_{\pm 0.072}$ 42.97 $\textbf{26.30}_{\pm 4.10}$ $\textbf{0.717}_{\pm 0.067}$ $\textbf{0.175}_{\pm 0.107}$ 64.19 DPS $\underline{22.73}_{\pm 6.07}$ $\underline{0.591}_{\pm 0.141}$ $0.264_{\pm 0.156}$ $112.82$ $19.23_{\pm 2.52}$ $0.582_{\pm 0.082}$ $0.503_{\pm 0.106}$ $146.23$ RED-diff $22.16_{\pm 3.41}$ $0.512_{\pm 0.083}$ $\underline{0.258}_{\pm 0.089}$ $\underline{108.32}$ $\underline{22.03}_{\pm 5.90}$ $\underline{0.601}_{\pm 0.094}$ $\underline{0.274}_{\pm 0.198}$ $\underline{113.48}$ Latent LatentDAPS(ours) $\textbf{25.94}_{\pm 2.87}$ $\textbf{0.751}_{\pm 0.056}$ $0.223_{\pm 0.080}$ $74.83$ $23.64_{\pm 4.10}$ $0.609_{\pm 0.053}$ $0.269_{\pm 0.099}$ $93.51$ ReSample $25.65_{\pm 3.57}$ $0.732_{\pm 0.059}$ $\textbf{0.182}_{\pm 0.085}$ 67.22 $\textbf{25.11}_{\pm 4.21}$ $\textbf{0.633}_{\pm 0.049}$ $\textbf{0.198}_{\pm 0.089}$ 87.66